# Project page: https://xiaohanlei.github.io/projects/GaussNav/

### v2

@@ Abstract
In embodied vision, Instance ImageGoal Navigation (IIN) requires an agent to locate a specific object depicted in a goal image within an unexplored environment. The primary difficulty of IIN stems from the necessity of recognizing the target object across varying viewpoints and rejecting potential distractors. Existing map-based navigation methods largely adopt the representation form of Bird’s Eye View (BEV) maps, which, however, lack the representation of detailed textures in a scene. To address the above issues, we propose a new Gaussian Splatting Navigation (abbreviated as GaussNav) framework for IIN task, which constructs a novel map representation based on 3D Gaussian Splatting (3DGS) [15]. The proposed framework enables the agent to not only memorize the geometry and semantic information of the scene, but also retain the textural features of objects. Our GaussNav framework demonstrates a significant leap in performance, evidenced by an increase in Success weighted by Path Length (SPL) from 0.252 to 0.578 on the challenging Habitat-Matterport 3D (HM3D) dataset. Our code will be made publicly available.
Keywords:
Embodied Visual Navigation 3D Gaussian Splatting
1 Introduction
@@
Figure 1: Instance ImageGoal Navigation (IIN) requires agent to navigate to the object instance depicted in the goal image, while distinguishing it from other visually similar instances of the same class.
As an emerging research topic, Embodied Artificial Intelligence (EAI) has been drawing an increasing interest in the community. This field demands that agents are capable of exploration, learning, reasoning, and interaction within the physical world. A critical question then arises: how can an embodied agent comprehend its world? Recent years have witnessed rapid advancements in radian field technologies, exemplified by the development of Neural Radiance Fields (NeRF) [27] and 3D Gaussian Splatting (3DGS) [15], among others. This leads to the subsequent inquiry: how can we leverage these radian field technologies to facilitate embodied tasks for agents?
To address this question, we focus on the embodied visual navigation [7, 18, 44, 22]. Traditional navigation has been studied over the past years, but with the advent of deep learning technologies, there is a renewed interest in navigating agents and also understanding the semantic intricacies of the 3D environment. For instance, consider a scenario where an agent is presented with a randomly captured image of a specific object, the agent’s task is to navigate to this object within a certain time budget. To distinguish the target from different angle of views, the agent has to understand the object’s geometry, its semantic attributes, and its spatial relationship to the environment. This task, named as Instance ImageGoal Navigation (IIN) was introduced by Krantz et al. [19] for embodied visual navigation, as depicted in Fig. 1.
@@
Recently, map-based visual navigation methods [4, 5, 17, 29, 1, 25] have focused on constructing explicit episodic Bird-Eye-View (BEV) maps, achieving notable success in established visual navigation benchmarks. These well-designed explicit map representations are memory-efficient, storing essential information such as 2D geometry and semantics, and can be directly utilized to calculate the agent’s subsequent actions. However, this simple 2D BEV map representation is not free from its limitations. For instance, it lacks the capacity to retain 3D geometrical information about the environment, rendering it ineffective for navigating cross-floor scenarios. Additionally, BEV maps are unable to preserve the intricate textures and detailed features in a scene. To address the above issues in existing map-based visual navigation methods, we formulate a new Gaussian Splatting Navigation framework, i.e., GaussNav, for IIN task.
@@
Our GaussNav framework consists of three stages, Sub-gaussians Division, Semantic Gaussian Construction and Gaussian Navigation. First, to address the optimization difficulties of 3DGS in large scenes on computationally constrained devices, we propose dividing gaussians into corresponding sub-gaussians and optimizing each of them individually. Second, we introduce a new form of map representation, i.e., Semantic Gaussian. By leveraging Semantic Segmentation algorithms, we assign semantic labels to each gaussian. We then cluster gaussians with their semantic labels and 3D positions, segmenting objects in the scene into different instances under various semantic categories. This representation is capable of preserving not only the 3D geometry of the scene and the semantic labels of each gaussian but also the texture details of the scene, thereby enabling novel view synthesis. Third, we render descriptive images for object instances, matching them with the goal image to effectively locate the target object. Upon determining the predicted goal object’s position, we can efficiently transform our Semantic Gaussian into grid map and employ path planning algorithms.
With our proposed method, GaussNav, the difficult IIN task can be transformed into a more tractable PointGoal Navigation task, consequently yielding more promising results. On the challenging Habitat-Matterport 3D dataset (HM3D) [41], we achieve $0.725$ success and $0.578$ SPL. This represents a significant improvement over the previous state-of-the-art work [21], which reported success of $0.702$ and SPL of $0.252$.
2 Related Work
We briefly discuss related work on ImageGoal, Instance ImageGoal Navigation (IIN) and 3DGS below.
ImageGoal Navigation. ImageGoal Navigation requires the agent to navigate to the location where the goal image is captured, matching the position and orientation accurately. The goal image is taken with parameters identical to those of the agent’s RGB-D camera, including agent’s height, field-of-view, and image size. To tackle this challenge problem, researchers adopt various strategies. Some optimize RL policy [1, 8, 39, 40] that directly maps observations to actions. Others focus on constructing maps [20, 6, 30, 16], or on developing carefully hand-crafted matching algorithm [35]. While these methods yield promising results in ImageGoal Navigation, they do not entirely address the unique demands of the IIN task.
Instance ImageGoal Navigation. The IIN task introduces distinct challenges compared to the ImageGoal Navigation. First, the goal image in IIN must feature a specific object instance, rather than meaningless large white walls in the ImageGoal Navigation. Second, the parameters used to capture the goal image, do not necessarily match those of the agent’s camera. Therefore, to succeed in IIN, an agent must be adept at identifying the target object among numerous candidates of the same class with goal object and recognizing it from various viewpoints. To address these challenges, Krantz et al. [17] develop a general pipeline for aligning the same object from different angle of views, Bono et al. [3] present an end-to-end approach in the IIN task, and Lei et al. [21] propose a method that mimics human behaviour for verifying objects at a distance. Existing methods focus partly on designing more sophisticated modules [17, 21] and partly on pre-training on pretext tasks [3]. Our approach differs in that we concentrate on designing a new map representation. Through this novel form of map, we can better establish the connection between target descriptions and target locations, thereby facilitating visual navigation.
@@ 2 Related Work
3D Gaussian Splatting. 3DGS, proposed by Bernhard et al. [15], models a 3D scene through anisotropic 3D gaussians. This scene representation demonstrates promising results on both rendering speed and quality. Leveraging these advantages, a growing body of research begins to explore various innovations, including deformable or dynamic gaussians [24, 36, 42], advancements in mesh extraction and physics simulation [10, 38, 11], as well as applications in Simultaneous Localization and Mapping (SLAM) [14, 26, 43]. This surge of interest in the capabilities of 3DGS leads us to consider its potential for embodied visual navigation. Given that the gaussian representation inherently encapsulates explicit scene geometry and the parameters required for rendering, we posit that 3DGS can significantly enhance embodied visual navigation. The explicit nature of the gaussian representation provides a rich, condensed form of environmental data, which can be effectively utilized to inform and guide autonomous agents in navigation.
3 Methods
3.1 Preliminaries
@@
In the IIN task, at the start of a new episode $e$, the agent is given a goal image $I_{g}$ that features a specified object instance $o$. The agent’s task is to navigate to this instance within the environment. At each timestep $t$, the agent acquires observations which include an RGB image $I_{t}$, a depth image $D_{t}$, and sensor pose reading $P_{t}$. Utilizing this information, the agent must decide upon and execute an action $a_{t}$. When calling stop action, the episode is success only if the agent is within a certain range of the goal object.
@@
We propose a modular framework called Gaussian Splatting for Visual Navigation (GaussNav) to tackle the IIN task. In a new environment, the Instance ImageGoal Navitation in our proposed framework consists of three stages: Sub-gaussians Division, as shown in Fig. 2, Semantic Gaussian Construction, as depicted in Fig. 2(a), and Gaussian Navigation, illustrated in Fig. 2(b). Initially, during the first episode within a unknown environment, the agent employs frontier exploration to explore the environment and collect observations. We first divide the observations into subsets and use our proposed Semantic Gaussian to reconstruct each of the sub-scene. Subsequently, in the following episodes, the agent leverages the Semantic Gaussian to grounding the object instance $o$ depicted in the goal image and navigate to it. This process effectively transforms the IIN task into a more manageable PointGoal Navigation task.
3.2 Sub-gaussians Division
Figure 2: Sub-gaussians Division. When initialized in an unkonwn environment, agent first employs frontier exploration to explore the environment. The agent will next partition the collected observations into distinct subsets for constructing sub-gaussians, in preparation for the subsequent Semantic Gaussian Construction.
@@ 3.2 Sub-gaussians Division
In the first episode of an unexplored environment, the agent simultaneously maintains two types of maps, an exploration map and an obstacle map. The exploration map delineates the regions of the environment that have been explored, while the obstacle map marks the obstacles in the scene. By detecting the contours of the exploration map and excluding areas in the obstacle map, the agent set the closest frontier point as waypoint to facilitate exploration. By the applying frontier-based exploration strategy, the agent collects observations of the entire environment. As depcited in Fig. 2, we divide observations along the trajectory into distinct subsets. Then each subset is used to construct Semantic Gaussian in the next stage of Semantic Gaussian Construction.
3.3 Semantic Gaussian Construction
(a) Semantic Gaussian Construction.
(b) Gaussian Navigation.
Figure 3: The last two stages of our GaussNav framework. (a) Semantic Gaussian Construction. At timestep $t$, the pipeline updates the gaussians from $t-1$ through densification and updating, which involves a comparison between the rendered RGB and depth images against the current input training views. Concurrently, semantic labels are assigned to the densified gaussians using the segmented images. Finally, the gaussians are refined through differentiable rendering. (b) Gaussian Navigation. Our approach begins with the classification of the goal image using pre-constructed Semantic Gaussian. Upon determining the predicted class, we generate descriptive images around instances belonging to that class. These images are then matched with the target object to identify and ground the goal instance. Utilizing the map and the established goal, the agent employs path planning to compute the sequence of actions.
We formulate our Semantic Gaussian Construction problem as, given a set of camera poses and RGB-D images, construct a semantic gaussian map. The semantic gaussian map can represent scene’s geometry, feature scene’s semantic information and render high-fidelity RGB-D image given an arbitrary camera pose. At any timestep $t$, the agent’s RGB observation $I_{t}$ is segmented into $C_{t}$ using Mask-RCNN [12]. Specifically, $C_{t}$ is a $6\times H\times W$ mask, where 6 is the number of possible goal categories in the IIN task. At $t=0$, the Semantic Gaussian is initialized with a semantically labeled point cloud based on the observed RGB $I_{0}$, depth image $D_{0}$, and segmentation result $C_{0}$. Subsequently, the gaussians are initialized and optimized based on the point cloud. When $t>0$, Gaussian Densification is performed by comparing the rendered results at position $P_{t}$ with the ground truth, followed by an update to the semantic Gaussian through differentiable rendering, as depicted in Fig. 2(a).
Gaussian Map Representation. The Semantic Gaussian of the scene is represented as a set of semantically labeled gaussians. The representation of gaussians is simplified compared to 3DGS [15] by using only view-independent color and constraining gaussians to be isotropic. Consequently, each gaussian is characterized by a minimal set of eight parameters: a triplet for the RGB color vector $\mathbf{c}$, a triplet delineating the centroid $\bm{\mu}\in\mathbb{R}^{3}$, a scalar representing the radius $r$, a scalar quantifying the opacity $o\in[0,1]$, and a scalar representing the category label $l$. Each gaussian influences a point in 3D space $\mathbf{x}\in\mathbb{R}^{3}$ according to the standard unnormalized Gaussian equation weighted by its opacity:

### v3

# GaussNav: Gaussian Splatting for Visual Navigation

@@ Abstract
In embodied vision, Instance ImageGoal Navigation (IIN) requires an agent to locate a specific object depicted in a goal image within an unexplored environment. The primary challenge of IIN arises from the need to recognize the target object across varying viewpoints while ignoring potential distractors. Existing map-based navigation methods typically use Bird’s Eye View (BEV) maps, which lack detailed texture representation of a scene. Consequently, while BEV maps are effective for semantic-level visual navigation, they are struggling for instance-level tasks. To this end, we propose a new framework for IIN, Gaussian Splatting for Visual Navigation (GaussNav), which constructs a novel map representation based on 3D Gaussian Splatting (3DGS). The GaussNav framework enables the agent to memorize both the geometry and semantic information of the scene, as well as retain the textural features of objects. By matching renderings of similar objects with the target, the agent can accurately identify, ground, and navigate to the specified object. Our GaussNav framework demonstrates a significant performance improvement, with Success weighted by Path Length (SPL) increasing from 0.347 to 0.578 on the challenging Habitat-Matterport 3D (HM3D) dataset. The source code is publicly available at the link: https://github.com/XiaohanLei/GaussNav.
Index Terms:
Embodied Visual Navigation, 3D Gaussian Splatting.
I Introduction
@@
Figure 1: Illustration of Instance ImageGoal Navigation (IIN), which requires agent to navigate to the object instance depicted in the goal image, while distinguishing it from other visually similar instances.
Embodied visual navigation is an emerging computer vision problem where an agent uses visual sensing to actively interact with the world and perform navigation tasks [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Recent years have witnessed substantial progress in embodied visual navigation, fueled by the availability of large-scale photo-realistic 3D scene datasets [13, 14, 15] and fast simulators for embodied navigation [16, 17, 18]. These advancements have enabled researchers to develop and test navigation algorithms in controlled environments that closely mimic real-world conditions.
In embodied visual navigation, one critical question is “how can we describe the goal target”. One common setting for tasking an agent is to give it natural language instructions, such as “check if my laptop is on the chair”. However, this setting becomes confusing when there are multiple chairs in the house. To overcome this challenge, Krantz et al.[19] propose Instance ImageGoal Navigation (IIN) [20, 21, 22]. In IIN, an agent is presented with an image of a specific object, and its goal is to navigate to the specific object within the least time budget. The goal image is not expected to match the sensor specification or embodiment of the navigating agent, as described in Figure 1. To accomplish the task, the agent needs to distinguish the target object from different angle of views and ignore potential distractors. This is a challenging task as it involves semantic reasoning, geometry understanding and instance-aware matching.
@@
To address the above issue, previous methods [23, 24, 20, 25, 26, 27, 22] introduce 2D Semantic Bird-Eye-View (BEV) map to tackle this problem. These well-designed explicit map representations are memory-efficient, storing essential information such as 2D geometry and semantics, and can be directly utilized to calculate the agent’s subsequent actions. However, this simple 2D BEV map representation lacks the capacity to retain 3D geometrical information about the environment, rendering it ineffective for navigating cross-floor scenarios. Additionally, BEV maps are unable to preserve the instance-aware features in a scene, which can be crucial for distinguishing between multiple objects of the same class or for tasks requiring fine-grained object interaction.
To avoid the limitation of BEV maps, we propose a new Gaussian Splatting Navigation framework, i.e., GaussNav, for IIN task. Our GaussNav is inspired by the recent advancements in 3D Vision technologies, including Neural Radiance Fields (NeRF) [28] and 3D Gaussian Splatting (3DGS) [29]. These technologies have demonstrated substantial progress in novel view synthesis (NVS) and 3D scene understanding. Although these technologies were not initially developed for navigation tasks, they offer considerable potential for application in this domain. To this end, we develop the Semantic Gaussian map representation, which integrates the representation of geometry, semantics and instance-aware features, and can be directly used for visual navigation.
@@
Our GaussNav framework consists of three stages, including Frontier Exploration, Semantic Gaussian Construction and Gaussian Navigation. First, the agent employs Frontier Exploration to collect observations of the unknown environment. Second, the collected observations are used to construct Semantic Gaussian. By leveraging semantic segmentation algorithms [30, 31], we assign semantic labels to each Gaussian. We then cluster Gaussians with their semantic labels and 3D positions, segmenting objects in the scene into different instances under various semantic categories. This representation is capable of preserving not only the 3D geometry of the scene and the semantic labels of each Gaussian, but also the texture details of the scene, thereby enabling NVS. Third, we render descriptive images for object instances, matching them with the goal image to effectively locate the target object. Upon determining the predicted goal object’s position, we can efficiently transform our Semantic Gaussian into grid map and employ path planning algorithms to accomplish the navigation.
To the best of our knowledge, we are the first to introduce 3DGS [29] to embodied visual navigation. In this work, we unify the map representation of geometry, semantics and instance-aware features for visual navigation. Our framework can directly ground the target object with a single goal image input and guide the agent towards it without any additional exploration or verification [22]. Our framework designs are beneficial for effective and efficient visual navigation. We evaluate our method’s on both efficacy and efficiency and establish new state-of-the-art records on the challenging Habitat-Matterport 3D dataset (HM3D) [14].
II Related Work
We briefly discuss related work on differentiable rendering, followed by a broad overview of embodied visual navigation, and finally, we focus on the work most relevant to us: Instance ImageGoal Navigation (IIN).
Differentiable Rendering. To achieve photo-realistic scene capture, differentiable volumetric rendering has gained prominence with the introduction of Neural Radiance Fields (NeRF) [28]. NeRF utilizes a single Multi-Layer Perceptron (MLP) to represent a scene, performing volume rendering by marching along pixel rays and querying the MLP for opacity and color. Due to the inherent differentiability of volume rendering, the MLP representation is optimized to minimize rendering loss using multi-view information, resulting in high-quality novel view synthesis (NVS). The primary limitation of NeRF is its slow training speed. Recent advancements have addressed this issue by incorporating explicit volume structures, such as multi-resolution voxel grids [32, 33, 34] and hash functions [35], to enhance performance.
@@ II Related Work
In contrast to NeRF, 3DGS [29] employs differentiable rasterization. Unlike ray marching, which iterates along pixel rays, 3DGS iterates over the primitives to be rasterized, similar to conventional graphics rasterization. By leveraging the natural sparsity of a 3D scene, 3DGS achieves an expressive representation capable of capturing high-fidelity 3D scenes while offering significantly faster rendering. Comprehensive review of the developments in 3D Gaussian Splatting [36] highlights the method’s versatility and applications across various domains. Leveraging these advantages, a growing body of research begins to explore various innovations, including deformable or dynamic Gaussians [37, 38, 39], advancements in mesh extraction and physics simulation [40, 41, 42], as well as applications in Simultaneous Localization and Mapping (SLAM) [43, 44, 45]. This surge of interest in the capabilities of 3DGS leads us to consider its potential for embodied visual navigation. Given that the Gaussian representation inherently encapsulates explicit scene geometry and the parameters required for rendering, we posit that 3DGS can enhance decision-making in map-based visual navigation. The explicit nature of the Gaussian representation provides a rich, condensed form of environmental data, which can be effectively utilized to inform and guide autonomous agents in navigation.
Embodied Visual Navigation. Embodied Visual navigation includes several topics: ObjectGoal Navigation (ObjectNav), Multi ObjectGoal Navigation (MultiON), ImageGoal Navigaiton (ImageNav), and Instance ImageGoal Navigation (IIN). ObjectNav [46, 24, 47, 48, 49] requires an agent to navigate to any instance of a specified object category within the environment. MultiON [50, 51, 52, 53], on the other hand, tasks the agent to sequentially navigate to a series of objects. ImageNav [26, 54, 55, 56, 57, 58, 59, 60, 61] involves navigating to the camera pose from which a target image is captured. In contrast, IIN [20, 21, 22] requires navigating to the specific instance captured by the camera in the target image. Collectively, these navigation tasks span a spectrum from semantic-level navigation in ObjectNav and MultiON to fine-grained instance-level navigation in ImageNav and IIN, comprehensively capturing the problem space of embodied visual navigation.
Numerous approaches to solving embodied visual navigation utilize deep reinforcement learning (DRL) to develop end-to-end policies that map egocentric vision to action [26, 46, 56, 62]. However, acquiring skills related to visual scene understanding, semantic exploration, and long-term memory are challenging in an end-to-end framework. Consequently, these methods often incorporate a combination of careful reward shaping [54], pre-training routines [56], and advanced memory modules [59, 63, 64]. In contrast to end-to-end DRL, alternative approaches decompose the problem into sub-tasks that can be optimized in a supervised manner. These sub-tasks include graph prediction via topological SLAM [58], graph-based distance learning [65, 66], and camera pose estimation for last-mile navigation [61]. Chaplot et al.[24] decompose the embodied visual navigation task into exploration, object detection, and local navigation. Building on this, CLIP on Wheels (CoW) [67] utilizes a similar decomposition strategy, focusing on exploration and object localization to handle an open-set object vocabulary. Modular approaches also show promise for effective simulation-to-reality transfer (Sim2Real). Gervet et al.[68] conduct a Sim2Real transfer of both modular and end-to-end systems, demonstrating that modularity effectively mitigates the visual Sim2Real gap that impairs the performance of end-to-end policies.
Modular approaches typically represent the environment using a map and utilize this map to acquire the knowledge necessary for navigation within that environment. Bird’s-Eye View (BEV) maps [69, 24, 20, 22] are a form of metric map that project the entire scene from an overhead perspective, representing the occupancy status, exploration state, and semantic categories of specific areas. This map representation encodes information about each region into a grid, achieving accurate spatial awareness. In contrast, topological maps [58, 60, 57] focus more on describing the spatial relationships between different areas in a scene. This approach allows agents to navigate based on the connectivity of spaces rather than relying solely on geometric coordinates. To represent a scene more meticulously, 3D-aware maps have been proposed [49, 70, 71, 72]. Incorporating 3D consistency can enhance perception in navigation. Building upon 3D maps, we further advance by adopting 3D Gaussian Splatting (3DGS) [29] as a novel map representation. This enables the map to synthesize appearance views of specific object instances, which we have demonstrated to be effective in the Instance ImageGoal Navigation task.
Instance ImageGoal Navigation. The IIN task introduces distinct challenges compared to the ImageGoal Navigation. First, in IIN, the goal image must depict a specific object instance. In contrast, ImageGoal Navigation may use randomly captured photos that could include insignificant elements, such as large white walls. Second, the camera parameters used to capture the goal image do not necessarily match those of the agent’s camera. Therefore, to succeed in IIN, an agent must be adept at identifying the target object among numerous candidates of the same class with goal object and recognizing it from various viewpoints. To address the above challenges, Krantz et al.[20] develop a general pipeline for aligning the same object from different angle of views. Bono et al.[21] present an end-to-end approach in the IIN task, while Lei et al.[22] propose a method that mimics human behaviour for verifying objects at a distance. Existing methods focus partly on designing sophisticated modules [20, 22] and partly on pre-training on pretext tasks [21]. Our approach differs in that we concentrate on designing a new map representation. Through this novel form of map, we can better establish the connection between target description and target locations, thereby facilitating visual navigation.
III Methods
Figure 2: Framework Overview. In the first episode of a scene, the agent uses frontier exploration to gather observations of the unknown environment, constructing a Semantic Gaussian. In subsequent episodes, the pre-constructed Semantic Gaussian is utilized by Gaussian Navigation to ground the goal object and guide the agent towards it.
Figure 3: Exploration Map and Obstacle Map.
III-A Overview
@@
In the IIN task, at the start of a new episode $e$, the agent is given a goal image $I_{g}$ that features a specified object instance $O_{g}$. The agent’s goal is to navigate to the referred instance $O_{g}$ within the environment. At each timestep $t$, the agent acquires observations which include an RGB image $I_{t}$, a depth image $D_{t}$, and sensor pose reading $P_{t}$. Utilizing this information, the agent must decide upon and execute an action $a_{t}$. When calling stop action, the episode is considered successful only if the agent is within a certain range of the goal object.
@@
To accomplish the IIN task, we propose a modular framework called Gaussian Splatting for Visual Navigation (GaussNav), as depicted in Figure 2. In a new environment, the Instance ImageGoal Navitation in our proposed framework consists of three stages: Frontier Exploration, Semantic Gaussian Construction, as depicted in Figure 4, and Gaussian Navigation, illustrated in Figure 5. Initially, during the first episode within an unknown environment, the agent employs frontier exploration to explore the environment and collect observations. We then use our proposed Semantic Gaussian to reconstruct the scene. Subsequently, in the following episodes, the agent leverages the Semantic Gaussian to ground the object instance $o$ depicted in the goal image and navigate to it. This process effectively transforms the IIN task into a more manageable PointGoal Navigation task.
III-B Frontier Exploration
@@
In the first episode of an unexplored environment, the agent simultaneously maintains two types of maps, an exploration map and an obstacle map, as illustrated in Figure 3. The exploration map delineates the regions of the environment that have been explored, while the obstacle map marks the obstacles in the scene. By detecting the contours of the exploration map and excluding areas in the obstacle map, the agent sets the closest frontier point as a waypoint to facilitate exploration. The frontier-based exploration strategy is a well-established approach in robotics and autonomous navigation [73, 74]. It involves identifying the boundaries between the explored and unexplored regions of the environment, known as frontiers. The agent then selects the nearest accessible frontier point as its next target for exploration. This decision is based on the distance to the frontier, the navigability of the path, and the potential information gain from exploring that frontier [75]. By iteratively exploring the nearest frontiers, the agent efficiently covers the entire environment while avoiding obstacles and previously explored areas. Through the application of this frontier-based exploration strategy, the agent collects observations of the entire environment.
III-C Semantic Gaussian Construction
Our Semantic Gaussian is represented by a group of Gaussians, with each Gaussian characterized by a minimal set of nine parameters: a triplet for the RGB color vector $\mathbf{c}$, a triplet delineating the centroid $\bm{\mu}\in\mathbb{R}^{3}$, a scalar representing the radius $r$, a scalar quantifying the opacity $o\in[0,1]$, and a scalar representing the category label $l$. Different from original 3DGS [29], we simplify the representation of Gaussian by using only view-independent color and constraining the Gaussian to be isotropic. This simplification enhances computational efficiency while reducing memory requirements. For a comprehensive understanding of 3DGS [29], we highly recommend readers consulting the original paper [29]. Our Semantic Gaussian Construction can be described as an iterative process comprising two alternating steps: Gaussian Densification and Semantic Gaussian Updating, as depicted in Figure 4. Gaussian Densification initializes new Gaussians in the Semantic Gaussian at each new incoming frame, while Semantic Gaussian Updating refines the parameters of each Gaussian through Differentiable Rendering.
Differentiable Rendering. 3DGS [29] renders an RGB image as follows: given a collection of 3D Gaussians and camera pose, first sort all Gaussians from front to back. RGB images can then be efficiently rendered by alpha-compositing the splatted 2D projection of each Gaussian in order in pixel space. The rendered color of pixel ${\bf p}=(u,v)$ can be written as:
where $f_{i}(\bf{p})$ is computed as follows: